# Leveraging Content and Acoustic Representations for Speech Emotion Recognition

### v2

@@ Abstract
Speech emotion recognition (SER), the task of identifying the expression of emotion from spoken content, is challenging due to the difficulty in extracting representations that capture emotional attributes from speech. The scarcity of labeled datasets further complicates the challenge where large models are prone to over-fitting. In this paper, we propose CARE (Content and Acoustic Representations of Emotions), where we design a dual encoding scheme which emphasizes semantic and acoustic factors of speech. While the semantic encoder is trained using distillation from utterance-level text representations, the acoustic encoder is trained to predict low-level frame-wise features of the speech signal. The proposed dual encoding scheme is a base-sized model trained only on unsupervised raw speech. With a simple light-weight classification model trained on the downstream task, we show that the CARE embeddings provide effective emotion recognition on a variety of datasets. We compare the proposal with several other self-supervised models as well as recent large-language model based approaches. In these evaluations, the proposed CARE is shown to be the best performing model based on average performance across diverse datasets. We also conduct several ablation studies to analyze the importance of various design choices.
@@
Speech Emotion Recognition (SER) focuses on detecting the speaker’s emotional state from the audio signal. Recognizing emotions in speech has significant applications across diverse fields, including human-computer interaction [1], social media analysis [2], customer service call centers [3], and mental health monitoring systems [4]. However, despite considerable progress, SER continues to pose challenges due to the complexity of human emotions and the inherent difficulties in effectively capturing these emotions from limited labeled datasets.
@@
Traditionally, SER systems have relied on various acoustic properties of speech signals to interpret emotional content. Lieberman et al.[5] were among the first to emphasize the role of pitch contour in emotion analysis. Additional acoustic features, including energy, intensity, and speaking rate, were later recognized as important indicators of emotional class [6]. The features identified through the Interspeech para-linguistic challenges were rich in emotional properties while being inherently high-dimensional [7, 8]. Eyben et al. [9] introduced a minimalist feature set to address this dimensionality issue partially.
In recent years, the field of speech emotion recognition (SER) has shifted to the adoption of deep learning methods. Current network architectures in SER commonly include convolutional neural networks (CNN)[10, 11], long short-term memory (LSTM) networks[12], and transformer models [13]. While these models perform well on the specific datasets they are trained on, they often struggle to generalize across diverse datasets. Due to the difficulty of training individual models for each dataset, self-supervised learning (SSL) models have emerged as a promising solution.
@@
Self-supervised learning (SSL) models, trained on large volumes of unlabeled data, have shown impressive generalization capabilities across a range of downstream tasks. Notable examples include wav2vec 2.0 [14], HuBERT [15], and WavLM [16]. These models are engineered to capture speech patterns in a manner similar to how models like BERT [17] capture language patterns. Although trained on neutral speech data, these models have demonstrated encouraging results in emotion recognition tasks [18, 19]. The emotion recognition performance may be additionally enhanced by training these models with tailored self-supervised objective functions that are emotion aware. Two recent examples are Vesper [20] and emotion2vec [21]. Both methods share a common strategy of utilizing a student-teacher framework for distilling knowledge from an existing model, such as WavLM (in the case of Vesper) or data2vec (in the case of emotion2vec), while training on emotional datasets. However, emotion is shaped not only by these features but also by the semantic content of the speech signal [22]. For instance, identifying emotions from text transcripts is often more effective than interpreting them from raw audio [19]. The integration of speech content during the pre-training phase of SER models remains an under-explored yet promising area of research.
@@
In this work, we introduce a self-supervised model for speech emotion recognition (SER) called Content and Acoustic Representations of Emotions (CARE). To the best of our knowledge, our approach is the first effort to pre-train a self-supervised model that integrates both semantic and acoustic components of speech. CARE leverages a dual encoding framework for processing speech signals: a semantic encoder, which aligns speech representations with sentence-level transcripts by adapting a pre-trained text representation model using CNNs, and a non-semantic encoder, which aligns speech representations with low-level acoustic features from the PASE+ model [23]. The outputs of both encoders are combined, and a lightweight classification head is then trained to perform emotion recognition across multiple downstream datasets. The key contributions of our work are as follows:-
@@
Proposing a novel self-supervised model for speech emotion recognition (SER) consisting of dual encoders: a semantic encoder and an acoustic encoder. Both encoders, being trained using distillation-based losses, enable parameter efficient learning.
@@
Developing an adaptation strategy for aligning pre-trained text models with speech inputs using convolutional adapters.
@@
Experimenting on eight benchmark speech datasets with diverse SER tasks, showcasing the effectiveness of CARE.
@@ II-A Audio Feature Extraction for SER
Recently, deep learning-based representations have gained popularity as low-level acoustic features. Notable examples include the SincNet architecture by Ravanelli et al.[24] and interpretable Gaussian filters by Agrawal et al.[25]. The LEAF front-end [26], was utilized by Dutta et al. [11] for speech emotion classification. Typically, these models require end-to-end training of both feature extractors and classifiers to generate meaningful speech representations for downstream tasks. In contrast, the proposed CARE architecture is a self-supervised model designed to generalize across diverse datasets without needing to fine-tune the pre-trained weights.
@@ II-B Self-supervision for SER
Among the general purpose speech SSL models that were proposed over the years, WavLM [16], was shown to outperform other models such as HuBERT [15] and wav2vec2.0 [14] for emotion recognition. Vesper [20] used a modified masking strategy to emphasize high pitch/energy regions of speech—known indicators of emotion—and derived targets for these masked regions from a WavLM teacher model. This approach allowed to learn a smaller student model with enhanced performance for SER. A similar strategy was employed by Ma et al. in emotion2vec [21], which utilized a pre-trained data2vec model as the teacher. Emotion2vec [21] also learns a global embedding to enhance SER performance. In contrast, the proposed CARE model integrates semantic content along with acoustic features to perform emotion recognition. A distillation loss from a text model enables the semantic encoding in a parameter efficient way.
II-C Speech-text Aligned Representations
@@
With the increasing prominence of large language models (LLMs), recent approaches have integrated speech encoders with LLMs. Notably, the SALMONN model by Tang et al.[30] introduced an audio encoder consisting of Whisper model and a music encoder along with the LLaMA language model [31]. Hu et al.[32] proposed WavLLM, combining Whisper and WavLM encoders with the LLaMA model. These LLM-based approaches harness aligned speech-text representations and enable prompt-based applications. However, their substantial model sizes (e.g., B parameters for SALMONN [30]) present significant computational demands for both training and inference.
In contrast, the CARE model achieves superior performance on various downstream datasets with a much smaller size of M parameters.
@@ III-A1 RoBERTa
One of the significant contributions in creating a text representation model was proposed by Devlin et al. [17]. This architecture, called bidirectional encoder representations from transformer (BERT), was trained with two objectives on a corpus of textual data, namely, predicting words masked out in a sentence and to predict whether two sentences semantically follow each other (referred to as next sentence prediction). Liu et. al [33] trained this architecture on a larger corpus of textual data without the next sentence prediction task. This pre-trained model, known as robust optimized BERT approach (RoBERTa), was shown to outperform BERT in a number of downstream tasks.
@@ III-B CARE Model
Having discussed the need for both semantic and acoustic supervision for our model in Sec. I, we propose a dual encoding scheme for modeling speech. These dual encoders — the semantic and acoustic encoders — process the speech signal through distinct supervisory signals suited to their respective objectives. The chosen supervision for each encoder is detailed as follows:
@@
An intuitive target for training a model to understand the semantics (or content) of speech would be the corresponding text transcripts. However, obtaining annotated transcripts for large emotional speech pre-training datasets is resource-intensive. Consequently, we assume the pre-training dataset lacks ground-truth text transcripts (even though we have access to it in our case) to develop a methodology that remains applicable even to datasets without human-generated transcriptions. In such a scenario, pre-trained automatic speech recognition (ASR) systems offer an alternative for generating these transcripts, and we use the open-source ASR model Whisper-large-v3 [34] for this task.
Before delving into further details, we wish to discuss a small caveat here. ASR systems have been shown in various studies to exhibit higher word error rates (WER) on emotional speech compared to neutral speech datasets [22]. However, Li et al. [22] show that since the datasets considered in their study consist of many short utterances, the ASR model does not have enough context to generate the correct transcript. This leads us to choose a dataset with podcast recordings as our pre-training corpus. Podcast recordings provide sufficiently long utterances that improve transcript accuracy and offer a broad content variety suitable for pre-training the semantic encoder. Specifically for our pre-training dataset, we observe a WER of which is not substantially high.
Given these generated transcripts, the next question is how best to train the semantic encoder. Since the semantic encoder’s purpose is to align the speech signal with its content to facilitate emotion recognition, an ASR-style alignment loss could be applied. However, a sentence level representation for text is more appropriate for the task of emotion recognition as established by Fan et al. [35]. Therefore, we extract contextual word-level embeddings from the transcripts using a pre-trained RoBERTa model[33] and mean-pool these embeddings to obtain a single feature vector representing the entire transcript. These utterance-level embeddings serve as the supervisory signal, or “teacher,” for the semantic encoder in our CARE model. We denote this utterance level embeddings by .
@@
While utterance level embeddings serve as the targets for the semantic encoder, frame level descriptors serve as targets for the acoustic encoder. An attentive reader might enquire as to why such a choice is made. Note that mean-pooled speech representations can capture a wide range of properties—such as speaker identity, accent, and language—beyond emotion alone [36]. Moreover, emotion in speech is often contained in the change of different parameters such as pitch, rhythm or energy from one frame to another [6]. Thus a frame-level target is chosen for the acoustic encoder.
A straightforward approach for these frame level acoustic targets would involve masking parts of the speech signal and reconstructing them, an approach adopted in both Vesper [20] and emotion2vec [21], the studies most closely related to ours. However, random masking is less effective for emotion recognition than selectively masking high-energy or high-pitch regions, as demonstrated by Chen et al. [20]. This highlights the role of low-level speech descriptors—such as filter-bank energies and pitch—as they are rich with emotional cues.
Based on these insights, we choose to predict PASE+ features, which encompass filter-bank energies, pitch, and other low-level descriptors essential for capturing emotion. Specifically, we use frame-level PASE+ features with dimensions as targets for the acoustic encoder in our CARE model. These features are down-sampled by a factor of , producing target descriptors at a frequency of Hz. We denote the acoustic targets from the PASE+ model by .
@@ III-B1 Model Architecture
The speech signal is first processed through a series of convolutional layers designed to produce frame representations every ms. These are followed by a stack of six transformer layers, forming the common encoder that serves both the acoustic and semantic encoder pathways in our model.
Recall that the target for the semantic encoder is obtained from the mean-pooled pre-trained RoBERTa representations for the generated text transcripts. The semantic encoder is thus expected to align the speech representations with its corresponding generated transcript. This encoder consists of six transformer layers which are initialized with the weights from a pre-trained text representation model. Being trained with textual data, the transformer layers in the semantic encoder do not generalize to speech representations. To address this, we propose a novel adaptation strategy by introducing two 1D-convolutional blocks—one placed before and one after each transformer layer.
The first block adjusts the speech representations from the common encoder to align them with the internal representations expected by the text-based model. The second block refines these representations post-transformer processing. Additionally, following established practice for processing speech in text models [37, 38, 30], the time-resolution of the speech sequence is reduced before processing by the transformer layers in the semantic encoder. Specifically, the convolutional block preceding each transformer layer down-samples the sequence length by a factor of three, while the block following it up-samples it by the same factor. Each convolutional block consists of a single convolutional layer, with a kernel size of , and input and output channels set to in order to match the dimension of the pre-trained transformer layers. While adaptation of speech SSL models with convolution layers has been explored in prior works [39, 40], adapting pre-trained text models for speech tasks, using convolutional adapters, is explored for the first time in this work. Importantly, the transformer layers themselves are not updated during training. Finally, the semantic encoder’s output representations are average-pooled to produce an utterance-level representation.
@@ III-B2 Loss
where denotes the acoustic target.
@@ III-B3 Inference
In this setup, the only learnable parameters for the downstream tasks are the weights for the convex combination and those of the lightweight classification head; the weights of the common, acoustic, and semantic encoders remain frozen during supervised training.
@@ IV-A Pre-training
The usage of podcast recordings for pre-training has already been motivated in Sec. III. Raw audio data from the MSP-PODCAST corpus [42] is used for the task of pre-training. A total of speech-turns amounting to hours of emotional speech data are used. Out of the total number of samples, of the data is randomly chosen as the training set while the remaining serves as the validation set. The Whisper-large-v3 model111https://huggingface.co/openai/whisper-large-v3 is used for generating the transcripts (the WER observed is ), while the pre-trained RoBERTa model222https://huggingface.co/FacebookAI/roberta-base is used for encoding the transcripts. The common encoder is initialized with first layers of the WavLM-base model333https://huggingface.co/microsoft/wavlm-base, while the acoustic encoder is initialized with the last layers of the same. The convolutional feature extractor is also initialized from the WavLM-base model. The transformer layers of the semantic encoder are initialized with the weights of the last layers of a pre-trained RoBERTa base model, while the convolutional adapters are randomly initialized.
@@ IV-C Loss and Evaluation Metrics for Downstream Tasks
Since these datasets are categorical in nature, we use the cross entropy loss for training the downstream model weights (the convex combination weights and the lightweight classification head parameters). For testing, we use the weighted F1-score as the evaluation metric as many of the datasets are class-imbalanced in nature (See Table I) . Denoting the F1 score of a class with samples, by , the weighted F1-score is defined as:
@@ IV-D1 Pre-training
During pre-training, all the speech utterances from MSP-PODCAST are padded or randomly cropped to a duration of seconds. The model is trained with a learning rate of - and a batch size of with AdamW [56] as the optimizer. The model is trained for a total of steps and the best model parameters based on validation set performance are chosen for evaluation of downstream datasets.

### v3

@@ Abstract
Speech emotion recognition (SER), the task of identifying the expression of emotion from spoken content, is challenging due to the difficulty in extracting representations that capture emotional attributes. The scarcity of labeled datasets further complicates the challenge where large models are prone to over-fitting. In this paper, we propose CARE (Content and Acoustic Representations of Emotions), where we design a dual encoding scheme which emphasizes semantic and acoustic factors of speech. While the semantic encoder is trained using distillation from utterance-level text representations, the acoustic encoder is trained to predict low-level frame-wise features of the speech signal. The proposed dual encoding scheme is a base-sized model trained only on unsupervised raw speech. With a simple light-weight classification model trained on the downstream task, we show that the CARE embeddings provide effective emotion recognition on a variety of datasets. We compare the proposal with several other self-supervised models as well as recent large-language model based approaches. In these evaluations, the proposed CARE is shown to be the best performing model based on average performance across diverse datasets. We also conduct several ablation studies to analyze the importance of various design choices.
@@
Speech Emotion Recognition (SER) focuses on detecting the speaker’s emotional state from the audio signal. Recognizing emotions in speech has significant applications across diverse fields, including human-computer interaction [1], social media analysis [2], customer service call centers [3], and mental health monitoring systems [4]. However, despite considerable progress, SER continues to pose challenges due to the complexity of human emotions and the inherent difficulties in effectively capturing them from limited labeled datasets.
@@
Traditionally, SER systems have relied on various acoustic properties of speech signals. Lieberman et al.[5] emphasize the role of pitch contour in emotion analysis, while additional acoustic features, including energy, intensity, and speaking rate, were recognized as indicators of emotional class [6]. The features identified through the Interspeech para-linguistic challenges were rich in emotional properties while being high-dimensional [7, 8]. Eyben et al. [9] introduced a minimalist feature set to address this dimensionality issue.
@@
In recent years, the network architectures in SER commonly include convolutional neural networks (CNN)[10, 11], long short-term memory (LSTM) networks[12], and transformer models [13]. While these models perform well on the specific datasets, they often struggle to generalize across diverse datasets. In such settings, self-supervised learning (SSL) models have emerged as a promising solution. Notable examples of SSL approaches include wav2vec 2.0 [14], HuBERT [15], and WavLM [16]. These models are engineered to capture speech patterns similar to textual models like BERT [17]. Although trained on neutral speech data, these models have demonstrated encouraging results in emotion recognition tasks [18, 19]. The emotion recognition performance may be further enhanced by training these models with emotion-aware self-supervised objectives. Two recent examples are Vesper [20] and emotion2vec [21]. However, emotion in speech is also shaped by its semantic content [22]. For instance, identifying emotions from text transcripts is often more effective than interpreting them from raw audio [19]. The integration of speech content during the pre-training phase of SER models remains an under-explored yet promising area of research.
@@
In this work, we introduce a self-supervised model for speech emotion recognition (SER) called Content and Acoustic Representations of Emotions (CARE). To the best of our knowledge, our approach is the first effort to pre-train a self-supervised model that integrates both semantic and acoustic components of speech. CARE leverages a dual encoding framework for processing speech signals: a semantic encoder, which aligns speech representations with sentence-level transcripts, and a non-semantic encoder, which aligns speech representations with low-level acoustic features from the PASE+ model [23]. The outputs of both encoders are combined, and a lightweight classification head is then trained to perform emotion recognition. The key contributions are :-
@@
Proposing a novel self-supervised model for speech emotion recognition (SER) consisting of dual encoders: a semantic encoder and an acoustic encoder.
@@
Developing an adaptation strategy for aligning pre-trained text models with speech inputs by convolutional adapters.
@@
Experimenting on benchmark speech datasets with diverse tasks, showcasing the effectiveness of CARE.
@@ II-A Audio Feature Extraction for SER
Recently, deep learning-based representations have gained popularity as low-level acoustic features. Notable examples include the SincNet architecture by Ravanelli et al.[24] and interpretable Gaussian filters by Agrawal et al.[25]. The LEAF front-end [26], was utilized by Dutta et al. [11] for speech emotion classification. Typically, these models require end-to-end training of both feature extractors and classifiers. In contrast, the proposed CARE architecture is a self-supervised model designed to generalize across diverse datasets.
@@ II-B Self-supervision for SER
Among the general purpose speech SSL models that were proposed over the years, WavLM [16], was shown to outperform other models such as HuBERT [15] and wav2vec2.0 [14] for emotion recognition. Vesper [20] used a modified masking strategy to emphasize high pitch/energy regions of speech—known indicators of emotion—and derived targets for these masked regions from a WavLM teacher model. A similar strategy was employed by Ma et al. in emotion2vec [21], which utilized a pre-trained data2vec model as the teacher. Emotion2vec [21] also learns a global embedding to enhance SER performance. In contrast, the proposed CARE model integrates semantic content along with acoustic features.
II-C Multimodal Emotion Recognition
The use of speech signals alongside text transcripts for multimodal emotion recognition has been explored in several prior works [11, 29, 30]. These approaches typically involve separate modeling of the two modalities, followed by a fusion stage. In contrast, CARE is designed to model the semantic and acoustic properties of speech with the uni-modal input.
II-D Speech-text Aligned Representations
@@
With the increasing prominence of large language models (LLMs), recent approaches have integrated speech encoders with LLMs. Notably, the SALMONN model by Tang et al.[32] introduced an audio encoder consisting of Whisper model and a music encoder along with the LLaMA language model [33]. Hu et al.[34] proposed WavLLM, combining Whisper and WavLM encoders with the LLaMA model. These LLM-based approaches harness aligned speech-text representations, enabling prompt-based applications. However, their substantial model sizes (e.g., B parameters for SALMONN) present significant computational demands for both training and inference.
In contrast, CARE achieves superior performance on various downstream datasets with a much smaller size of M parameters.
@@ III-A1 RoBERTa
One of the significant contributions in creating a text representation model was proposed by Devlin et al. [17]. Liu et. al [35] trained this architecture on a larger corpus of textual data without the next sentence prediction task. This pre-trained model, known as robust optimized BERT approach (RoBERTa), was shown to outperform BERT in a number of downstream tasks.
@@ III-B CARE Model
We propose a dual encoding scheme (semantic and acoustic encoders) to process the speech signal through distinct supervisory signals suited to their respective objectives. The chosen supervision for each encoder is detailed as follows:
@@
We do not assume the availability of ground-truth text transcripts for the pre-training data. In such a scenario, pre-trained automatic speech recognition (ASR) systems (Whisper-large-v3 [36]) offer an alternative for generating these transcripts.
Typically, ASR systems have been shown to exhibit higher word error rates (WER) on emotional speech compared to neutral speech datasets [22].
Podcast recordings, on the other hand, provide sufficiently long context and offer a broad content variety suitable for pre-training the semantic encoder. Specifically, we observe a WER of , which may be reasonable for SER tasks.
Since the semantic encoder’s purpose is to align the speech signal with its content to facilitate emotion recognition, an ASR-style alignment loss could be applied. However, a sentence-level representation for text is more appropriate for the task of emotion recognition as established by Fan et al. [37]. Therefore, we extract contextual word-level embeddings from the transcripts using a pre-trained RoBERTa model [35] and mean-pool these embeddings to obtain a single feature vector representing the entire transcript. These utterance-level embeddings serve as the supervisory signal, or “teacher”, for the semantic encoder in our CARE model. We denote these utterance-level embeddings by .
@@
In prior works, mean-pooled representations have shown to encode characteristics like speaker identity, accent, and language [38]. However, we speculate that
emotion in speech is often contained in fine-grained acoustic attributes such as pitch, rhythm, and their modulations [6]. Thus, a frame-level target is chosen for the acoustic encoder.
A direct approach for the frame level acoustic targets would involve masking parts of the speech signal and reconstructing them. However, prior works show that random masking is less effective for emotion recognition than selectively masking high-energy or high-pitch regions, as demonstrated by Chen et al. [20]. Based on these observations, we choose to predict PASE+ features, which encompass filter-bank energies, pitch, and other low-level descriptors essential for capturing emotion. Specifically, we use frame-level PASE+ features with dimensions as targets for the acoustic encoder in our CARE model. These features are down-sampled by a factor of , producing target descriptors at a frequency of Hz. We denote the acoustic targets from the PASE+ model by .
@@ III-B1 Model Architecture
The speech signal is first processed through a series of convolutional layers designed to produce frame representations every ms. These are followed by a stack of six transformer layers, forming the common encoder that serves both the acoustic and semantic encoder pathways in the proposed model.
The semantic encoder is designed to align the speech representations with its corresponding generated transcript. This encoder consists of six transformer layers which are initialized with the weights from a pre-trained text representation model. Being trained with textual data, the transformer layers in the semantic encoder do not generalize to speech representations. To address this, we propose a novel adaptation strategy by introducing two 1D-convolutional blocks—one placed before and one after each transformer layer.
The first block adjusts the speech representations from the common encoder to align them with the internal representations expected by the text-based model. The second block refines these representations post-transformer processing. Additionally, following established practice for processing speech in text models [39, 40, 32], the time resolution of the speech sequence is reduced before processing by the transformer layers in the semantic encoder. Specifically, the convolutional block preceding each transformer layer down-samples the sequence length by a factor of three, while the block following it up-samples it by the same factor. Each convolutional block consists of a single convolutional layer, with a kernel size of , and input and output channels set to in order to match the dimension of the pre-trained transformer layers. While adaptation of speech SSL models with convolution layers has been explored in prior works [41, 42], adapting pre-trained text models for speech tasks, using convolutional adapters, is explored for the first time in this work. Importantly, the transformer layers themselves are not updated during training. Finally, the semantic encoder’s output representations are average-pooled to produce an utterance-level representation.
@@ III-B2 Loss
where denotes the acoustic target. The total loss during pre-training is given by
where is decided based on the validation performance.
@@ III-B3 Inference
In this setup, the only learnable parameters for the downstream tasks are the weights for the convex combination and those of the lightweight classification head.
@@ IV-A Pre-training
The MSP-PODCAST corpus [44] is used for the task of pre-training. A total of samples amounting to hours of emotional speech data are used. Out of these, of the data is randomly chosen as the training set while the remaining serves as the validation set. The Whisper-large-v3 model is used for generating the transcripts (the WER observed is ), while the pre-trained RoBERTa model is used for encoding the transcripts. The common encoder is initialized with first layers of the WavLM-base model, while the acoustic encoder is initialized with the last layers of the same. The convolutional feature extractor is also initialized from the WavLM-base model. The transformer layers of the semantic encoder are initialized with the weights of the last layers of a pre-trained RoBERTa base model, while the convolutional adapters are randomly initialized.
@@ IV-C Loss and Evaluation Metrics for Downstream Tasks
The cross-entropy loss is used for training the downstream model weights (the convex combination weights and the lightweight classification head parameters). For testing, we use the weighted F1-score as the evaluation metric as many of the datasets are class-imbalanced (Table I). Denoting the F1 score of class with samples, by , the weighted F1-score is
We also report the unweighted average recall (UAR) for all cases, which is the mean of the class-wise recall scores.
@@ IV-D1 Pre-training
During pre-training, all the speech utterances from MSP-PODCAST are padded or randomly cropped to a duration of seconds. The model is trained with a learning rate of - and a batch size of with AdamW [58] as the optimizer. The model is trained for a total of steps and the best model parameters based on validation set performance are chosen for evaluation of downstream datasets. We experiment with different values of (Eq. 3) to balance the two losses during pre-training. Setting results in degraded performance, while increasing it to does not yield any significant improvement over . Therefore, we fix for all the subsequent experiments.